10.6 g of anhydrous sodium carbonate combines with 18g of water to form hydrated sodium carbonate whose chemical formula is …. (Na=23, C=12, O=16, H=1)

1
Expert's answer
2022-01-20T04:08:02-0500

10.6+18/23+12+1=28.6/36=0.8